5. Calculate the RBC indices for patient D: RBC count 4.79 x 106/uL, Hgb 12.5 g/dl, Het 36.0 % . Provide the correct RBC indices and RBC description. 6. Calculate the RBC indices for patient D: RBC count 5.55 x 10s/uL, Hgb 9.8 g/dl, Hct 25.0 % . Provide the correct RBC indices and RBC description.
Answer #1

(5).

The three different red blood cell indices include,

  1. Mean corpuscular volume (MCV)
  2. Mean corpuscular hemoglobin (MCH)
  3. MCH concentration (MCHC)

The percentage of cells in the liquid is known as the “hematocrit (HCT),” it is also known as packed cell volume (PCV) in the blood. The decreased hematocrit value indicates that the blood cell count is decreased. White blood cells (WBCs) are a type of red blood cells.

MCV (Mean corpuscular volume) = %hematocrit (HCT) × 10/ Red blood cell (RBC) count (millions/mm3)

MCH indicates the average weight of hemoglobin in each RBC molecule.

MCH = Hemoglobin / RBC count

MCHC indicates the average weight of hemoglobin compared to the cell size

MCHC = Hemoglobin / haematocrit

Red cell distribution width refers to the degree of variation in red blood cells. The reference range is, 9 to 14.5

Thus, as per the given information,

MCV = (36 × 10) / 4.79 = 75.15 fL

The MCV is within the normal range of 60 to 77, so the cells are normocytic.

MCH = (12.5/100 mL) / 4.79*10^6   = 26 pg

The normal range of MCH is, 25 to 33 pg. So, the hemoglobin weight is less than the normal

MCHC = (12.5/100 mL) / 36.0% = 34.7%